a 20% ammonia solution and a 5% ammonia solution are to be mixed to obtain 50 gallons of an 8% solution. how much of each should be used fill out a tabel
Answer by checkley77(12844)

.20X+.05(50-X)=.08*50
.20X+2.5-.05X=4
.15X=4-2.5
,15X=1.5
X=1.5/.15
X=10 GAL. OF 20% SOLUTION IS NEEDED.
50-10=40 GAL. OF 5% SOLUTION IS NEEDED.
PROOF:
.20*10+.05*40=.08*50
2+2=4
4=4